# A Pairwise Comparison Relation-assisted Multi-objective Evolutionary Neural Architecture Search Method with Multi-population Mechanism

### v1

@@ IV-A Experimental Configurations
Architecture Top-1 Acc (%) Top-5 Acc (%) MAdds (M) Search Cost (GPU Days) Search Method
EfficientNet-B0 [12] 76.3 93.2 390 - manual
EfficientNet-B1 [12] 78.8 94.4 700 - manual
MobileNetV2 [45] 72.0 91.0 300 - manual
PC-DARTS [46] 75.8 92.7 597 3.8 GD
P-DARTS [47] 75.6 92.6 557 0.3 GD
SNAS [55] 72.7 90.8 522 1.5 GD
$\beta$-DARTS [56] 76.1 93.0 609 0.4 GD
NAP [57] 75.5 92.6 574 4 GD
Shapely-NAS [58] 76.1 - 582 4.2 GD
OnceForAll [43] 76.9 93.2 230 2 EA
FairNAS [52] 77.5 - 392 12 EA
NSGANetV1 [15] 76.9 93.0 292.5 12 EA
CARS [35] 75.2 92.5 591 0.4 EA
MUXNet [54] 76.6 93.2 318 11 EA
MixPath [59] 77.2 93.5 378 10.3 EA
SLE-NAS-B [60] 75.7 92.5 412 3 EA
NASNet-A [6] 74.0 91.6 564 1800 RL
MnasNet [19] 76.13 92.85 391 - RL
SMEMNAS-S(ours) 77.75 93.75 360 0.17 EA
SMEMNAS-M(ours) 78.16 93.92 468 0.17 EA
SMEMNAS-L(ours) 78.91 94.23 570 0.17 EA
Figure 5: Accuracy and number of multi-adds in millions on CIFAR-10. Models from multi-objective approaches are joined with lines. SMEM-NAS finds more accurate solutions with fewer complexity.
IV-B Results on Standard Datasets
In this section, we show the results of our method on benchmark datasets to validate its effectiveness. We conduct the search and train the final architectures separately on three datasets. This is different from existing NAS methods that mostly use a transfer learning setup. We think the transferred architectures maybe not optimal on another dataset. Then, in order to validate the effectiveness of the proposed method, we compare the non-dominated architectures obtained by SMEM-NAS with those obtained by other state-of-the-art NAS methods. The selected peer methods can be broadly divided into three categories: manually designed by human experts, EA-based, and non-EA-based (e.g., RL and GD). In addition, we give the searched results by SMEM-NAS without the multi-population mechanism.
Results on CIFAR-10: After completing the search using our method, we selected three promising architectures, and for ease of representation we named them SMEMNAS-S/M/L (sorted by MAdds). Fig. 5 and Table I show the results of SMEM-NAS and the comparison with other methods on CIFAR-10. Fig. 5 visualizes the dominant relationship between the network models in terms of accuracy and model complexity. It is obvious that our models always outperform most models or our models are comparable to them. Firstly, from the perspective of time, our method takes 1.25 GPU days and significantly reduces the search time compared with methods based on EA and RL. The surrogate model we used effectively accelerates the search process. Compared with the GD-based method, although there are some deficiencies in time, the architecture we searched has better performance. Our model achieves the accuracy of 98.13%, surpassing almost all peer competitors. Among them, our model is slightly lower than FairNAS-A, but has nearly 100M lower MAdds and the search time is much less than FairNAS. In general, SMEM-NAS is superior to or consistent with other models in different metrics.
Results on CIFAR-100: Similar to the settings on CIFAR-10, we also conducted experiments on CIFAR-100, and Fig. 6 and Table II show the results of our model compared to networks from other works. We also provide three models on this dataset, large, medium and small, in which the highest accuracy can be 87.95%. From Fig. 6, it is clear that our architectures dominates most methods in terms of accuracy and complexity. Table II shows more details. On this dataset, although our results are slightly inferior to EfficientNet, the hand-crafted model requires significant expertise and labour. Additionally, we can see that the architecture with the highest accuracy searched on the CIFAR-10 dataset transferred to CIFAR-100 does not perform optimally, achieving 87.16% accuracy.
@@ IV-B Results on Standard Datasets
Results on ImageNet: Different from the conclusion on the CIFAR datasets, there is no need for pre-training before searching on the ImageNet dataset. Table III shows the results of the proposed method on the ImageNet dataset, SMEM-NAS outperforms other methods and achieves the highest accuracy of 78.91% with the MAdds of 570M. In addition, our proposed surrogate model effectively reduces the search time, and only takes 4 hours to complete the search.
@@
Figure 9: Search with different surrogate models. The top row shows the results of the search using four different models.The horizontal and vertical axes are error and MAdds, respectively. The bottom row shows the correlation between the predicted accuracy ranking and the true ranking using $KTau$.
@@ IV-C Performance of Multi-population Mechanism
The multi-population mechanism is an important strategy to increase the diversity of the solutions while ensuring convergence in this study. In order to verify its effectiveness, further experiments are conducted on the CIFAR-100 dataset in this section. We separately conduct the architecture search with and without the multi-population mechanism, and the comparison between the two search processes is shown in Fig. 7. The figures show the changes of the pareto front and the distribution of new individuals generated during the evolutionary process. Fig. 7a shows the search process without the multi-population mechanism and Fig. 7b shows the search process using the multi-population mechanism. Fig. 7 shows the distribution of the architectures searched in the archive, which have been realistically evaluated. Specifically, in both figures, the blue dots represent individuals at the initial and early stages of the evolutionary process, and the corresponding blue line is the pareto front at the early stage. In Fig. 7a, it is obvious that except for the randomly initialized individuals, the new individuals generated in the later evolution are all almost distributed in the same solution region, and there is no obvious change in the pareto front. However, in Fig. 7b, we can see that the new individuals generated are uniformly distributed unlike without the multi-population mechanism. The pareto front is also further updated in the excellent direction (lower error rate, smaller model size) and more promising individuals are explored. Evidently, the multi-population mechanism is better able to ensure the diversity of solutions and prevent them from falling into the local optima. Additionally, we show changes of the pareto front during the search process by using our method in Fig. 8. As we can see that the pareto front has stabilized in the early stage of the search, and is constantly changing and updating with the increase of iterations. In general, that is the same as our expected effect, accelerating the convergence speed while maintaining the diversity.
@@ IV-D Ablation Study
Comparison of classification models: In order to construct more accurate surrogate models, we consider four classification models: SVM, random forest (RF), k-nearest neighbors (KNN), and multi-layer perceptron (MLP). We perform four groups of experiments on the CIFAR-100 dataset. For more intuitive comparisons, we compute the correlation of 300 fully trained and evaluated architectures during the search process with each surrogate model. We choose the Kendall’s Tau coefficient ($KTau$) to show the correlation between their true rankings and predicted rankings. Fig. 9 shows the searched results using different models and the corresponding $KTau$. Architectures existing in the archive are used to train surrogate models. After that, new $K$ architectures are retained in each generation and used for testing. The figure shows the results of the last generation. From the $KTau$ and the distribution of individuals in the search process, it can be seen that SVM is the best classifier with very limited training samples. The new individuals generated during the evolution when using KNN and RF as surrogate models are worse and non-uniformly distributed, which may be caused by the inaccurate prediction at the early stage. In addition, the performance of MLP is similar to SVM as shown in the figure. But, SVM can take less time for training and achieve better results than MLP with the limited samples. Therefore, we choose SVM as the surrogate model.
@@
Compared to traditional surrogate models, we conduct data augmentation, expanding from $n$ samples to $n(n-1)/2$, which greatly increases the training data. Increased training data can effectively improve the reliability of the surrogate model. Table IV shows the performance comparison of the proposed surrogate model with the traditional regression surrogate models. We use 300 samples for training models, and in each experiment each surrogate model is trained with the same data. We train three models with the proposed classification-based method (SVM, RF, and KNN), and we train another three models with the traditional regression-based method (MLP, Decision Tree, and AdaBoost).
Table IV highlights that while the training and prediction time of the proposed surrogate model surpasses that of traditional models, classification-based surrogate models demonstrate superior performance in terms of correlation coefficients. Among these, our SVM achieves the highest $KTau$ coefficient at 0.8113. Therefore, the proposed surrogate model can provide more reliable prediction results and is more capable of retaining and choosing the excellent individuals during the search. Moreover, although it takes more time to train the proposed surrogate model, the training and prediction only takes a few more minutes during the actual search, which is acceptable compared to the whole search time.
TABLE IV: Comparison with different surrogate models. We record the Spearman coefficients, $KTau$ coefficients between the predicted accuracy rankings and the true accuracy rankings. The results are the average of ten experiments carried out.
Type Model Trainning time (s) Prediction time (s) Kendall’s Tau Spearman’s Rho
@@
Figure 10: The hypervolume values are achieved with different size of initial samples.
TABLE V: Comparison with different sizes of initial samples to train surrogate model.
Kendall’s Tau Spearman’s Rho Initial size
@@
Hyper-parameters: The initial population is important for both the training of the surrogate model and the subsequent evolution. In order to achieve a better balance between search efficiency and search quality, we test the appropriate value of the initial population size, $N$, by conducting the following small-scale experiments on the CIFAR-10. We set different sizes of the initial population, i.e., $N=25$, $50$, $100$ and $150$. We record the effect of different population sizes on the hypervolume (HV) value and the comparison results are shown in Fig. 10. From the figure, it can be seen that HV achieves the highest value when the population size $N=100$. Although the HV obtained at $N=150$ and at $N=100$ are almost equivalent, 50% more individuals need to be evaluated when $N=150$, which is more time consuming. In addition, we test the prediction performance of the initial surrogate model as well. As seen in Table V, the reliability of the surrogate model is low if the initial population size is too small, which is another reason for the unfavourable search results at $N=25$ and $50$. Besides, if the surrogate model has low reliability, the search results are prone to be less stable. Therefore, after comprehensive consideration, we choose $N=100$ as the size of the initial population.
V Conclusion
@@
In this paper, we propose an evolutionary multi-objective algorithm by using a multi-population mechanism and a surrogate model based on pairwise comparison relations to perform neural architecture search. The proposed method utilizes a surrogate model to predict accuracy rankings of architectures without focusing on accuracy values. The proposed surrogate model makes full use of the limited training samples and can greatly reduce the search overhead. Besides, a multi-population mechanism is used to improve the diversity in multi-objective NAS, which not only increases the population diversity in the evolution process, but also accelerates the convergence of the algorithm. We have conducted a series of experiments on benchmark datasets to verify the effectiveness of SMEM-NAS. The final architectures found by SMEM-NAS outperform several existing search methods.
@@
In this work, our surrogate model only considers the accuracy and ignores other indicators of networks. Single-objective surrogate models are not effectively adapted to multi-objective search frameworks. Out next work plans to explore the multi-objective surrogate model to predict domination relations of architectures. Besides, it can be observed that the initial population is vital to the results of evolutionary algorithms. The proposed method simply uses random sampling to initialise the population, and the sampled architectures are non-uniformly distributed in the objective space, which may not be a proper approach for subsequent evolutionary exploration. Therefore, we intend to investigate efficient sampling methods to initialize the population by considering a given problem’s features.

### v2

@@ IV-B Results on Standard Datasets
Architecture Top-1 Acc (%) Top-5 Acc (%) MAdds (M) Params (M) Search Cost (GPU Days) Device Year Search Method
EfficientNet-B0 [13] 76.3 93.2 390 5.3 - - 2019 manual
PA-DARTS [57] 75.3 92.25 - 5.2 0.4 RTX 3090 2024 GD
RelativeNAS [60] 75.12 92.3 563 5.05 0.4 GTX 1080Ti 2023 GD
$\beta$-DARTS [72] 76.1 93.0 609 5.5 0.4 - 2022 GD
NAP [73] 75.5 92.6 574 4.8 4 RTX 2080Ti 2022 GD
iDARTS [54] 75.3 92.3 568 5.1 1.9 Tesla V100 2023 GD
SQNAS [74] 73.8 91.4 540 - 0.25 Tesla V100 2024 GD
Shapely-NAS [75] 76.1 - 582 5.4 4.2 - 2022 GD
RandomNAS [76] 74.2 - - 5.9 0.5 - 2025 GD
SPNAS [66] 78.62 94.07 687 6.6 0.37 RTX 3090 2025 EA
OnceForAll [51] 76.9 93.2 230 - 2 Tesla V100 2020 EA
FairNAS [49] 77.5 - 392 5.9 12 - 2021 EA
EEEA-Net-C2 [63] 76.2 92.7 217 6.0 - - 2021 EA
NSGANetV2 [25] 77.5 93.5 225 6.1 1 - 2020 EA
MixPath [77] 77.2 93.5 378 5.1 10.3 Tesla V100 2023 EA
DPS [78] 75.6 92.7 578 5.3 0.06 RTX A6000 2023 EA
CENAS-C [79] 79.6 94.4 482 4.98 1.98 RTX 2080Ti 2024 EA
SLE-NAS-B [80] 75.7 92.5 412 4.5 3 RTX A6000 2023 EA
MnasNet [20] 76.7 93.3 403 5.2 4.5 64*TPUv2 2019 RL
BNAS [69] 74.3 92.2 - 3.9 0.19 - 2022 RL
DBNAS [70] 77.6 93.5 386 4.9 0.9 RTX 3090 2025 RL
SMEMNAS-S(ours) 77.75 93.75 360 5.9 0.17 RTX 3090 - EA
SMEMNAS-M(ours) 78.16 93.92 468 6.6 0.17 RTX 3090 - EA
SMEMNAS-L(ours) 78.91 94.23 570 7.0 0.17 RTX 3090 - EA
@@
Results on ImageNet: Different from the conclusion on the CIFAR datasets, there is no need for pre-training before searching on the ImageNet dataset. Table III shows the results of the proposed method and other comparative methods. SMEM-NAS outperforms most other methods and achieves the accuracy of 78.91% with the MAdds of 570M. In addition, our proposed surrogate model effectively reduces the search time, and only takes 4 hours to complete the search. It is worth mentioning that although the final accuracy achieved by our method is somewhat lower than that of CENAS-C, our search time and resource consumption are significantly lower than CENAS-C. We obtained the final results using only 10% of the search time with just one GPU.
@@
(a) Architectures distribution (main 0.7, vice 0.3)
(b) Architectures distribution (main 0.3, vice 0.7)
(c) Architectures distribution (both 0.5)
Figure 7: The comparison of results between different probabilities for using two populations on the ImageNet. Different probability values indicate the probability that the parents generating the next generation of offspring will come from the two populations. We plot the distribution of architectures in the archive during the search process in three figures. The horizontal and vertical axes are MAdds and error, respectively.
@@
Figure 10: Search with different surrogate models. The top row shows the results of the search using four different models. The horizontal and vertical axes are MAdds and error, respectively. The bottom row shows the correlation between the predicted accuracy ranking and the true ranking using $KTau$.
@@ IV-C Performance of Multi-population Mechanism
To further explore the advantages of different populations and design appropriate thresholds for the multi-population mechanism, we designed three experiments to determine the optimal use of the main and vice populations during evolution. We tested three configurations based on two populations: favoring the main population, favoring the secondary population, and using with equal probability. For each, we set a different probability for the two populations. When favoring the use of the main population, the probabilities of using the main and vice populations are 0.7 and 0.3. The values are reversed when favoring the vice population. In the last case, the probability of using both is 0.5. Fig. 7 shows the results of our experiments, and it can be seen that when the main population is used too much, the pareto front converges better, but architectures are more concentrated; when the vice population is used too much, architectures are more uniformly distributed, but the convergence effect is poorer; and when the main and vice populations are used with equal probability, the search effect is somewhere balanced. Therefore, in order to achieve better search results, both convergence and diversity should be considered. In the early stage of search, both populations are used for the convergence and exploration in the search space. In the middle stage, the main population should be used to achieve fast convergence, and in the late stage, we should use the vice population for more diversity exploration.
The multi-population mechanism is an important strategy to increase the diversity of the solutions while ensuring convergence in this study. In order to verify its effectiveness, more experiments were conducted on the CIFAR-100 dataset in this section. We separately conducted the architecture search with and without the multi-population mechanism, and the comparison between the two search processes is shown in Fig. 8. The figures show the changes of the pareto front and the distribution of new individuals generated during the evolutionary process. Fig. 8a shows the search process without the multi-population mechanism and Fig. 8b shows the search process using the multi-population mechanism. Fig. 8 shows the distributions of the architectures searched in the archive, which have been realistically evaluated. Specifically, in both figures, the blue dots represent individuals at the initial and early stages of the evolutionary process, and the corresponding blue line is the pareto front at the early stage. In Fig. 8a, it is obvious that except for the randomly initialized individuals, the new individuals generated in the later evolution are all almost distributed in the same solution region, and there is no obvious change in the pareto front. However, in Fig. 8b, we can see that the new individuals generated are uniformly distributed unlike without the multi-population mechanism. The pareto front is also further updated in the excellent direction (lower error rate, smaller model size) and more promising individuals are explored. Evidently, the multi-population mechanism is better able to ensure the diversity of solutions and prevent them from falling into the local optima. Additionally, we show changes of the pareto front during the search process by using our method in Fig. 9. As we can see that the pareto front has stabilized in the early stage of the search, and is constantly changing and updating with the increase of iterations. In general, that is the same as our expected effect, accelerating the convergence speed while maintaining the diversity.
@@ IV-D Ablation Study
Comparison of classification models: In order to construct more accurate surrogate models, we considered four classification models: SVM, random forest, k-nearest neighbors, and multi-layer perceptron. We performed four groups of experiments on the CIFAR-100 dataset. For more intuitive comparisons, we computed the correlation of 300 fully trained and evaluated architectures during the search process with each surrogate model. We choose the Kendall’s Tau coefficient ($KTau$) to show the correlation between their true rankings and predicted rankings. Fig. 10 shows the searched results using different models and the corresponding $KTau$. Architectures from the archive were used to train surrogate models. After that, new $K$ architectures were retained in each generation and used for testing. The figure shows the results of the last generation. From the $KTau$ and the distribution of individuals in the search process, it can be seen that SVM is the best classifier with very limited training samples. The new individuals generated during the evolution when using KNN and RF as surrogate models are worse and non-uniformly distributed, which may be caused by the inaccurate prediction at the early stage. In addition, the performance of MLP is similar to SVM as shown in the figure. But, SVM can take less time for training and achieve better results than MLP with the limited samples. Therefore, we choose SVM as the surrogate model.
@@
Compared to traditional surrogate models, we conducted data augmentation, expanding from $n$ samples to $n(n-1)/2$, which greatly increases the training data. Increased training data can effectively improve the reliability of the surrogate model. Table IV shows the performance comparison of the proposed surrogate model with the traditional regression surrogate models. We used 300 samples for training models, and in each experiment each surrogate model was trained with the same data. We trained three models with the proposed classification-based method (SVM, RF, and KNN), and we trained another three models with the traditional regression-based method (MLP, Decision Tree, and AdaBoost). Table IV highlights that while the training and prediction time of the proposed surrogate model surpasses that of traditional models, classification-based surrogate models demonstrate superior performance in terms of correlation coefficients. Among these, our SVM achieves the highest $KTau$ coefficient at 0.8113. Therefore, the proposed surrogate model can provide more reliable prediction results and is more capable of retaining and choosing the excellent individuals during the search. Moreover, although it takes more time to train the proposed surrogate model, the training and prediction only take a few more minutes during the actual search, which is acceptable compared to the whole search time.
TABLE IV: Comparison with different surrogate models. We record the Spearman coefficients ($\rho$), Kendall’s Tau coefficients ($KTau$) between the predicted accuracy rankings and the true accuracy rankings. The results are the average of ten experiments carried out.
Type Model
Trainning
time (s)
Prediction
time (s)
$KTau$ $\rho$
@@
TABLE V: The hypervolume values are achieved with different size of initial samples.
HV Initial size
0.8313 25
0.8339 50
0.8363 100
0.8360 150
TABLE VI: Comparison with different sizes of initial samples to train surrogate model.
$KTau$ $\rho$ Initial size
@@
Hyper-parameters: The initial population is important for both the training of the surrogate model and the subsequent evolution. In order to achieve a better balance between search efficiency and search quality, we tested the appropriate value of the initial population size, $N$, by conducting the following small-scale experiments on the CIFAR-10. We set different sizes of the initial population, i.e., $N=25$, $50$, $100$ and $150$. We recorded the effect of different population sizes on the hypervolume (HV) value and the comparison results are shown in Table. VI. From the table, it can be seen that HV achieves the highest value when the population size $N=100$. Although the HV obtained at $N=150$ and at $N=100$ are almost equivalent, 50% more individuals need to be evaluated when $N=150$, which is more time consuming. In addition, we tested the prediction performance of the initial surrogate model as well. As seen in Table VI, the reliability of the surrogate model is low if the initial population size is too small, which is another reason for the unfavourable search results at $N=25$ and $50$. Besides, if the surrogate model has low reliability, the search results are prone to be less stable. Therefore, after comprehensive consideration, we choose $N=100$ as the size of the initial population.
V Conclusion and Future Work
V-A Conclusion
@@
In this paper, we propose an evolutionary multi-objective algorithm by using a multi-population mechanism and a surrogate model based on pairwise comparison relations to perform neural architecture search. The proposed method utilizes a surrogate model to predict accuracy rankings of architectures without focusing on accuracy values. The proposed surrogate model makes full use of the limited training samples and can greatly reduce the search overhead. Besides, a multi-population mechanism is used to improve the diversity in multi-objective NAS, which not only increases the population diversity in the evolution process, but also accelerates the convergence of the algorithm. We have conducted a series of experiments on benchmark datasets to verify the effectiveness of SMEM-NAS. The final architectures found by SMEM-NAS outperform most existing search methods.
V-B Limitations and Future Work
@@
In this work, our surrogate model only considers the accuracy and ignores other indicators of networks. Single-objective surrogate models are not effectively adapted to multi-objective search frameworks. Our next work plans to explore the multi-objective surrogate model to predict domination relations of architectures. Besides, it can be observed that the initial population is vital to the results of evolutionary algorithms. The proposed method simply uses random sampling to initialise the population, and the sampled architectures are non-uniformly distributed in the objective space, which may not be a proper approach for subsequent evolutionary exploration. Therefore, we plan to investigate efficient sampling methods to initialize the population in future work by considering a given problem’s features. In addition, to test the performance of the model and its applicability in different computer vision tasks, we will also use a pretrained model on ImageNet as the backbone to test its performance on other tasks such as object detection and semantic segmentation.